# references.bib

### v2

@@ 5 Discussion
Supplementary Information SI 1 Additional benchmarking results
@@
Supplementary Table 6: Matthews correlation coefficient (MCC) for graph-level classification tasks from various domains, presented as mean $\pm$ standard deviation over 5 different runs, and including GCN and GIN. oom denotes out-of-memory errors, and n/a that the model is unavailable (e.g., node/edge features are not integers, which are required for Graphormer and TokenGT). The highest mean values are highlighted in bold.
@@
Supplementary Table 7: Accuracy for graph-level classification tasks from various domains, presented as mean $\pm$ standard deviation over 5 different runs, and including GCN and GIN. oom denotes out-of-memory errors, and n/a that the model is unavailable (e.g., node/edge features are not integers, which are required for Graphormer and TokenGT). The highest mean values are highlighted in bold.
@@
Supplementary Table 8: Matthews correlation coefficient (MCC) for 11 node-level classification tasks, presented as mean $\pm$ standard deviation over 5 different runs, including GIN and DropGIN. The number of nodes for the shortest path (sp) benchmarks is given in parentheses (based on randomly-generated ‘infected’ Erdős–Rényi (ER) graphs; SI 8 for details). Additional heterophily results are provided in Table 10. The highest mean values are highlighted in bold.
@@
Supplementary Table 9: Accuracy for 11 node-level classification tasks, presented as mean $\pm$ standard deviation over 5 different runs, including GIN and DropGIN. The number of nodes for the shortest path (sp) benchmarks is given in parentheses (based on randomly-generated ‘infected’ Erdős–Rényi (ER) graphs; SI 8 for details). Additional heterophily results are provided in Table 10. The highest mean values are highlighted in bold.
@@
Dataset ($\uparrow$) GraphSAGE GT
heterophily roman empire $0.833566667$$\pm$ $0.001800926$ $0.8379$$\pm$ $0.000608276$
amazon ratings $0.359366667$$\pm$ $0.004735328$ $0.336733333$$\pm$ $0.009238146$
minesweeper $0.652533333$$\pm$ $0.008450049$ $0.5743$$\pm$ $0.035674501$
tolokers $0.232333333$$\pm$ $0.018396286$ $0.3834$$\pm$ $0.011898319$
squirrel filtered $0.1381$$\pm$ $0.029983162$ $0.181866667$$\pm$ $0.007352777$
chameleon filtered $0.272966667$$\pm$ $0.031440473$ $0.2531$$\pm$ $0.039151373$
Supplementary Information SI 2 Additional time and memory results
@@
Supplementary Figure 1: The elapsed time for training for a single epoch on the dockstring dataset (in seconds), and the maximum allocated memory during this training epoch (GB).
Supplementary Information SI 3 Sourcing and licensing
Supplementary Information SI 4 Transfer learning setup
The transfer learning setup consists of randomly selected training, validation, and test sets of 25K, 5K, and respectively 10K molecules with GW calculations (from the total 133,885). This setup mimics the low amounts of high quality/fidelity data available in drug discovery and quantum simulations projects. In the transductive case, the entire dataset with DFT targets is used for pre-training (including the 10K test set compounds, but only with DFT-level measurements), while in the inductive setting the 10K set is completely excluded. Here, we perform transfer learning by pre-training a model on the DFT target for a fixed number of epochs (150) and then fine-tuning it on the subset of 25K GW calculations. In the transductive case, pre-training occurs on the full set of 133K DFT calculations, while in the inductive case the DFT test set values are removed (note that the evaluation is done on the test set GW measurements).
Supplementary Information SI 5 Limitations
@@
In terms of limitations, we highlight that the available libraries are not optimised for masking or custom attention patterns. This is most evident for very dense graphs (tens of thousands of edges or more). Memory efficient and Flash attention are available natively in PyTorch [NEURIPS2019_9015] starting from version 2.0, as well as in the xFormers library [xFormers2022]. More specifically, we have tested at least 5 different implementations of ESA: (1) leveraging the MultiheadAttention module from PyTorch, (2) leveraging the MultiHeadDispatch module from xFormers, (3) a manual implementation of multihead attention, relying on PyTorch’s scaled_dot_product_attention function, (4) a manual implementation of multihead attention, relying on xFormers’ memory_efficient_attention, and (5) a naive implementation. Options (1) - (4) can all make use of efficient and fast implementations. However, we have observed performance differences between the 4 implementations, as well as compared to a naive implementation. This behaviour is likely due to the different low-level kernel implementations. Moreover, Flash attention does not currently support custom attention masks as there is little interest for such functionality from a language modelling perspective.
@@
Although the masks can be computed efficiently during training, all frameworks require the last two dimensions of the input mask tensor to be of shape $(L_{n},L_{n})$ for nodes or $(L_{e},L_{e})$ for edges, effectively squaring the number of nodes or edges. However, the mask tensors are sparse and a sparse tensor alternative could greatly reduce the memory consumption for large and dense graphs. Such an option exists for the PyTorch native attention, but it is currently broken.
@@
Another possible optimisation would be to use nested (ragged) tensors to represent graphs, since padding is currently necessary to ensure identical dimensions for attention. A prototype nested tensor attention is available in PyTorch; however, not all the required operations are supported and converting between normal and nested tensors is slow.
@@
For all implementations, it is required that the mask tensor is repeated by the number of attention heads (e.g., 8 or 16). However, a notable bottleneck is encountered for the MultiheadAttention and MultiHeadDispatch variants described above, which require that the repeats happen in the batch dimension, i.e., requiring 3D mask tensors of shape $(B\times H,L,L)$, where $H$ is the number of heads. The other two efficient implementations require a 4D mask instead, i.e., $(B,H,L,L)$, where one can use PyTorch’s expand function instead of repeat. The expand alternative does not use any additional memory, while repeat requires ×$\times H$ memory. Note that it is not possible to reshape the 4D tensor created using expand without using additional memory.
@@
Finally, we noticed a limitation involving PyTorch’s nonzero() tensor method, which is required as part of the edge masking algorithm (Section 2). This is covered in detail in SI 9. Currently, the nonzero() method fails for very dense graphs. A fix would require an update to 64-bit integer limits.
Supplementary Information SI 6 Helper functions
@@
consecutive is a helper function that generates consecutive numbers starting from 0, with a length specified in its tensor argument as the difference between adjacent elements, and a second integer argument used for the last length computation, e.g., consecutive($[1,4,6]$, $10$) $=[0,1,2,0,1,0,1,2,3]$, and first_unique_index finds the first occurrence of each unique element in the tensor (sorted), e.g., first_unique_index($[3,2,3,4,2]$) $=[1,0,3]$. The implementations are available in our code base.
Supplementary Information SI 7 Experimental setup
@@
We follow a simple and universal experimental protocol to ensure that it is possible to compare the results of different methods and to evaluate a large number of datasets with high throughput. Below in Section SI 7.1 we described the grid search approach and the used search parameters. For the basic hyperparameters such as batch size and learning rate, we chose a number of reasonable hyperparameters and settings for all methods, regardless of their nature (GNN or attention-based). This includes the AdamW optimiser [loshchilov2018decoupled], more specifically the 8-bit version [dettmers2022bit], learning rate ($0.0001$), batch size ($128$), mixed precision training with the bfloat16 tensor format, early stopping with a patience of 30 epochs (100 for the very small datasets such as freesolv), and gradient clipping (set to the default value of $0.5$). Furthermore, we used a simple learning rate scheduler that halved the learning rate if no improvement was encountered for 15 epochs (half the early stopping patience). If these parameters led to out-of-memory errors, we attempted reducing the batch size by 2 until the error was fixed, or reducing the hidden dimension as a last resort.
@@
For a fair and comprehensive evaluation, we tune each algorithm for each dataset using a grid search approach and a selection of reasonable hyperparameters. These include the number of layers, the number of attention heads for GAT(v2) and graph transformers, dropout, and hidden dimensions. For graph-level GNNs, we evaluated configurations with 4 to 6 layers, hidden dimensions in $\{128,256,512\}$, the number of GAT heads in $\{8,16\}$, and GAT dropout in $\{0,0.2\}$. For node-level GNNs, we adapted the search since these tasks are more sensitive to overfitting, and used a number of layers in $\{1,2,4,6\}$, hidden dimensions in $\{64,128,256\}$, the same GAT heads and dropout settings, and dropout after each GNN layer in $\{0,0.2\}$. For Graphormer, TokenGT, and GraphGPS, we evaluate models with the number of layers in $\{4,6,8,10\}$, the number of attention heads in $\{4,8,16\}$, and hidden dimensions in $\{128,256,512\}$. For ESA, we generally focused on models with 6 to 10 layers, with SABs at the start and end and MABs in the middle. Hidden dimensions are selected from $\{256,512\}$ and the number of attention heads from $\{8,16,32\}$. We evaluate pre-LN and post-LN architectures, standard and gated MLPs, and different MLP hidden dimensions and number of MLP layers on a dataset-by-dataset basis. The best configuration is selected based on the validation loss and results are reported on the test set from 5 different runs. Based on recent reports on the performance of GNNs [luo2024classicgnnsstrongbaselines], we augmented all 6 GNN baselines with residual connections and normalisation layers for each graph convolutional layer. These strategies are not part of the original message passing specification but lead to substantial uplifts. For datasets with established splits, such as cifar10, mnist, or zinc, we use the available splits. For dockstring, only train and test splits are available, so we randomly extract 20,000 train molecules for validation. Otherwise, we generate our own splits with a train/validation/test ratio of 80%/10%/10%. All models are trained and evaluated using mixed-precision training with the bfloat16 tensor format.
SI 7.2 Metrics
@@
Given the scale of our evaluation, it is crucial to use an appropriate selection of performance metrics. To this end, we selected the metrics according to established and recent literature. For classification tasks, there is a growing consensus that Matthew’s correlation coefficient (MCC) is the preferred metric over alternatives such as accuracy, F-score, and the area under the receiver operating characteristic curve (AUROC or ROC-AUC) [Chicco2020, 9440903, Chicco2021, STOICA2024109511]. The AUROC in particular has been shown to be problematic [Chicco2023, 10.1145/1557019.1557021, 10.1145/3593013.3594100]. The MCC is an informative measure of a classifier’s performance as it summarises the four basic rates of a confusion matrix: sensitivity, specificity, precision, and negative predictive value [Chicco2023]. Similarly, the $\text{R}^{2}$ has been proven to be more informative than alternatives such as the mean absolute or squared errors for regression tasks [Chicco2021-bg]. Thus, our first choices for reporting results are the MCC and $\text{R}^{2}$, depending on the task. For comparison with leaderboard results and for specialised fields such as quantum mechanics, we also report comparable metrics (i.e., accuracy, mean absolute error, or root mean squared error).
SI 7.3 Baselines
@@
We include classic message passing baselines in the form of GCN, GAT, and GIN due to their recent resurgence against sophisticated graph transformers and their widespread use. We also include the improved GATv2 [brody2022how] to complement GAT, and PNA for being neglected in other works despite its remarkable empirical performance. We complete the message passing baselines by including DropGNN [wang2023mathscrnwl], a family of provably expressive GNNs that can solve tasks beyond 1-WL in an efficient manner by randomly dropping nodes. As in the original paper, we use GIN as the main underlying mechanism and label this technique DropGIN. Regarding transformer baselines, we select Graphormer, TokenGT, and GraphGPS, not only due to their widespread use, but also due to generally outperforming previous generation graph transformers such as SAN. This selection is balanced, in the sense that Graphormer and TokenGT are part of a class of algorithms that focuses on representing the graph structure through encodings and token identifiers, while GraphGPS relies on GNNs and is thus a hybrid approach.
Supplementary Information SI 8 Infected graph generation
@@
The infected graphs are generated using the InfectionDataset from PyTorch Geometric. We generated two Erdős–Rényi (ER) graphs with different sizes:
@@
Supplementary Information SI 9 Adaptations for Open Catalyst Project
@@
Extending a given model to work with 3D data is not trivial, as demonstrated by the follow-up paper dedicated to extending and benchmarking Graphormer on 3D molecular problems [shi2023benchmarking]. As described in that paper, for the Open Catalyst Project (OCP) data certain pre-processing steps are taken to ensure satisfactory performance. Concretely, a set of Gaussian basis functions is used to encode atomic distances, which are not used in their raw form. The idea of encoding raw quantities to achieve expressive and orthogonal representations has also been studied by Gasteiger et al. for DimeNet [gasteiger_dimenet_2020], taking things further and using Bessel functions. This idea is prevalent in the literature and shows some of the complications of working with 3D coordinates.

### v3

@@ Methods
[segment=0]
Acknowledgements
We are thankful for the PhD sponsorship awarded to D.B. and access to the Scientific Computing Platform within AstraZeneca.
Author Contributions
D.B., J.P.J., D.O., and P.L. were involved in conceptualising and reviewing the manuscript. D.B. designed and carried out experiments and produced the figures and tables. J.P.J., D.O., and P.L. jointly supervised the work.
Competing Interests
David Buterez has been supported by a fully funded PhD grant from AstraZeneca. J.P.J and D.O. are employees and shareholders at AstraZeneca. All other authors declare no competing interests.
{refsection}
[…]
An end-to-end attention-based approach for learning on graphs
Supplementary Information
Supplementary Information SI 1 ESA compared to GAT and SAN
GAT
Our masked self-attention module relies on masking within classical scaled dot product attention which comes with different projection matrices for keys, queries, and values. Additionally, we wrap the masked scaled dot product attention with layer normalisation and skip connections, the output of which is passed through an MLP. The latter is not done classically in GAT as part of its attention modules nor in the overall architecture. In contrast to GAT that operates over nodes, our masking operator is defined over sets of edges that are connected if they share a node in common. When it comes to the readout, GAT uses sum, mean or max and aggregates over nodes whereas our architecture relies on pooling by multi-head attention and aggregation over learnt representations of seed vectors inherent to that module. In terms of the encoder that is responsible for learning representations of set items, ours interleaves masked and self-attention layers vertically, whereas in GATs the stacked layers are entirely based on neighbourhood attention.
SAN
One of the main differences in relation to SAN’s architecture is the horizontal versus vertical combination of masked and self-attention. Although it is possible to set $\gamma=0$ for some layers and $\gamma=1$ for others, this avenue has never been explored in SAN and it would still result in a slightly different encoder as the query and key projections are different for the input graph and its complement. The empirical analysis in [kreuzer2021rethinking] also does not provide insights on vertical combinations of masked and self-attention, nor on the impact of attention pooling which is a differentiating factor and an important part of our architecture.
Supplementary Information SI 2 Node masking algorithm
The following algorithm illustrates the masking operation on nodes instead of edges.
1 from torch_geometric import unbatch_edge_index
2 function node_mask(batched_edge_index, batch_map, B, M)
       4 # batch_map maps nodes to graphs
       5 # B, M are the batch, respectively mask size
       6 mask $\leftarrow$ torch.full(size=(B, M, M), fill=False)
       9 edge_index $\leftarrow$ torch.cat(edge_index_list, dim=1)
[…]
Algorithm 1 Node masking in the PyTorch Geometric framework.
Supplementary Information SI 3 Additional benchmarking results
@@ Acknowledgements
Supplementary Table 2: The table reports the mean absolute error (MAE) and average precision (AP) for two long-range molecular benchmarks involving peptides. The molecular graph of a peptide is much larger than that of a small drug-like molecule and this makes the tasks well-suited for long-range benchmarking. All the results except the ones for ESA and TokenGT are extracted from [tonshoff2023where]. The number of layers for pept-struct, respectively pept-func is given as (/)($\cdot$/$\cdot$)( ⋅ / ⋅ ).
Dataset GCN (6/6) GIN (10/8) GPS (8/6) TokenGT (10/10) ESA (3/4)
pept-fn (AP $\uparrow$) $0.686$$\pm$ $0.005$ $0.6621$$\pm$ $0.0067$ $0.6534$$\pm$ $0.0091$ $0.6263465732336044$ $\pm$ $0.011736145935986075$ $0.7071010768413544$ $\pm$ $0.001466499593936558$
@@
Supplementary Table 7: Matthews correlation coefficient (MCC) for graph-level classification tasks from various domains, presented as mean $\pm$ standard deviation over 5 different runs, and including GCN and GIN. oom denotes out-of-memory errors, and n/a that the model is unavailable (e.g. node/edge features are not integers, which are required for Graphormer and TokenGT). The highest mean values are highlighted in bold.
@@
Supplementary Table 8: Accuracy for graph-level classification tasks from various domains, presented as mean $\pm$ standard deviation over 5 different runs, and including GCN and GIN. oom denotes out-of-memory errors, and n/a that the model is unavailable (e.g. node/edge features are not integers, which are required for Graphormer and TokenGT). The highest mean values are highlighted in bold.
@@
Supplementary Table 9: Matthews correlation coefficient (MCC) for 11 node-level classification tasks, presented as mean $\pm$ standard deviation over 5 different runs, including GIN and DropGIN. The number of nodes for the shortest path (sp) benchmarks is given in parentheses (based on randomly-generated infected Erdős–Rényi (ER) graphs; SI 11 for details). Additional heterophily results are provided in Table 11. The highest mean values are highlighted in bold.
@@
Supplementary Table 10: Accuracy for 11 node-level classification tasks, presented as mean $\pm$ standard deviation over 5 different runs, including GIN and DropGIN. The number of nodes for the shortest path (sp) benchmarks is given in parentheses (based on randomly-generated infected Erdős–Rényi (ER) graphs; SI 11 for details). Additional heterophily results are provided in Table 11. The highest mean values are highlighted in bold.
@@
Dataset ($\uparrow$) GraphSAGE GT Polynormer
heterophily roman empire $0.833566667$$\pm$ $0.001800926$ $0.8379$$\pm$ $0.000608276$ $0.9166992783546448$ $\pm$ $0.0018511145211087185$
amazon ratings $0.359366667$$\pm$ $0.004735328$ $0.336733333$$\pm$ $0.009238146$ $0.37747885584831237$ $\pm$ $0.0038234452239483767$
minesweeper $0.652533333$$\pm$ $0.008450049$ $0.5743$$\pm$ $0.035674501$ $0.7683401465415954$ $\pm$ $0.016896075342260147$
tolokers $0.232333333$$\pm$ $0.018396286$ $0.3834$$\pm$ $0.011898319$ $0.4098785161972046$ $\pm$ $0.017801612158082663$
squirrel filtered $0.1381$$\pm$ $0.029983162$ $0.181866667$$\pm$ $0.007352777$ $0.22971308728059134$ $\pm$ $0.020225244250503562$
chameleon filtered $0.272966667$$\pm$ $0.031440473$ $0.2531$$\pm$ $0.039151373$ $0.387384033203125$ $\pm$ $0.018208205032305608$
Supplementary Information SI 4 Additional Exphormer and Polynormer results
We evaluate Exphormer [shirzad2023exphormer], which extends GraphGPS by adding virtual global nodes and expander graphs, and Polynormer [deng2024polynormer], which is an interesting combination of graph attention and gating. We perform this additional evaluation on a smaller but representative selection of datasets: i) relatively large-scale datasets from computational chemistry and computer vision (zinc and mnist), ii) small-scale molecular datasets (from MoleculeNet), and iii) bioinformatics benchmarks (enzymes, proteins).
Polynormer was originally evaluated only on node classification tasks, whereas our main focus is on graph-level predictive performance. We have adapted this baseline to graph-level tasks by including a simple mean pooling layer, and then trained and evaluated tuned Polynormer models.
The results of our evaluation on these tasks, featuring both Polynormer and Exphormer, are provided in Tables 1 and 3. In line with expectations, Exphormer performs similarly to GraphGPS, being better on datasets such as mnist and proteins, but slightly underperforming on the small molecular datasets and zinc. Polynormer does not appear to be competitive on several of the larger graph-level datasets, but is otherwise close.
Supplementary Table 12: $\text{R}^{2}$ for MoleculeNet datasets, presented as mean $\pm$ standard deviation over $5$ runs. ESA and GPS results are reproduced from Table 1.
Dataset ($\uparrow$) GPS Exphormer Polynormer ESA
mn freesolv $0.8611225$$\pm$ $0.03203359047546806$ $0.88916$$\pm$ $0.014467404743076786$ $0.893307626247406$ $\pm$ $0.011083389197296983$ $0.9772682245913693$ $\pm$ $0.0008614222385922215$
lipo $0.7899625$$\pm$ $0.003816892551539786$ $0.7511966666666666$ $\pm$ $0.017414793519687027$ $0.7988451957702637$ $\pm$ $0.007303996469696557$ $0.8088115295794889$ $\pm$ $0.00737575964623688$
esol $0.910505$$\pm$ $0.002596627620587873$ $0.9074266666666667$ $\pm$ $0.010135587468584817$ $0.8952259023984274$ $\pm$ $0.006014030747248456$ $0.9440256449113372$ $\pm$ $0.0018959215104232857$
Supplementary Table 13: Matthews correlation coefficient (MCC) for classification datasets, presented as mean $\pm$ standard deviation over $5$ runs. ESA and GPS results are reproduced from Tables 3 and 3.
Dataset ($\uparrow$) GPS Exphormer Polynormer ESA
mn bbbp $0.70488$$\pm$ $0.043869435829515746$ $0.67454$$\pm$ $0.03035917159607623$ $0.7362912893295288$ $\pm$ $0.0270823847413422$ $0.8353317141532898$ $\pm$ $0.01387638635608058$
bace $0.6184459999999999$ $\pm$ $0.03174741917069795$ $0.60086$$\pm$ $0.017314135843292895$ $0.6116288006305695$ $\pm$ $0.04098625600729607$ $0.7209106683731079$ $\pm$ $0.019175842540950262$
bio enzymes $0.733944$$\pm$ $0.04514641008983991$ $0.7135766666666666$ $\pm$ $0.021507432048790298$ $0.627992570400238$ $\pm$ $0.0018268130018561908$ $0.7510042667388916$ $\pm$ $0.009278810475490992$
proteins $0.443208$$\pm$ $0.022000980341793867$ $0.5202666666666667$ $\pm$ $0.03218763168257853$ $0.5718311443924904$ $\pm$ $0.055292091764154096$ $0.5890437960624695$ $\pm$ $0.01668787306560422$
cv mnist $0.9795320000000001$ $\pm$ $0.0011450659369660959$ $0.98296$$\pm$ $0.0016178689687363318$ $0.9704129546880722$ $\pm$ $0.0016616749191475683$ $0.9863295316696167$ $\pm$ $0.00044475799669724145$
Supplementary Table 14: Mean absolute error (MAE) on the zinc dataset, presented as mean $\pm$ standard deviation over $5$ runs. ESA and GPS results are reproduced from Table 1.
Dataset ()($\downarrow$)( ↓ ) GPS Exphormer Polynormer ESA ESA (PE)
zinc $0.023842$$\pm$ $0.0069779309254248125$ $0.041458$$\pm$ $0.006516634100515389$ $0.10051885495583208$ $\pm$ $0.003341063410192649$ $0.02683207359883397$ $\pm$ $0.0007197092487851843$ $0.017331929457918168$ $\pm$ $0.0009253668251948539$
Supplementary Information SI 5 Additional time and memory results
Continuation of Figure 3, including time and memory results for the dockstring dataset.
@@
Supplementary Figure 1: (a). The elapsed time for training all evaluated models for a single epoch on the dockstring dataset (in seconds). Graphormer is not included as it runs out of memory. (b). The maximum allocated memory during the same training epoch (GB).
Supplementary Information SI 6 Memory scaling with sparse edge adjacency storage
It is worth noting the efficiency of optimised attention implementations. For a sequence length of 1 million ($10^{6}$), an efficient implementation calculates 1 trillion ($10^{12}$) attention scores without materialising the square attention matrix. To put this in the context of graph learning and ESA, we generated random Barabasi-Albert (BA) graphs using PyTorch Geometric, with varying numbers of edges that range from 200,000 to 2.6 million. We next train a model using a single self-attention layer (no masking) and representative model parameters (total model dimension 256, 8 attention heads, with MLPs and normalisation layers). Figure 2 (blue line) illustrates the peak memory use graphically, demonstrating the linear memory scaling (for the largest graphs, the kernel might be adjusting as the GPU memory limit of 40GB approaches).
We then measured and plotted the amount of memory required to store the sparse edge masking information, as opposed to a full dense matrix (green line in Figure 2). Clearly, this cost is relatively trivial even for graphs as large as 2.6 million edges.
We next hypothesised what an efficient implementation might look like by simply adding the sparse storage cost to the self-attention cost (orange line in Figure 2). Note that we are still assuming that full self-attention is being used; however, an efficient implementation could avoid hundreds of millions of unnecessary calculations.
Supplementary Figure 2: Memory usage for efficient attention versus the number of edges.
Supplementary Information SI 7 Limitations
@@
In terms of limitations, we highlight that the available libraries are not optimised for masking or custom attention patterns. This is most evident for very dense graphs (tens of thousands of edges or more). Memory efficient and Flash attention are natively available in PyTorch [NEURIPS2019_9015] starting from version 2.0, as well as in the xFormers library [xFormers2022]. More specifically, we have tested at least 5 different implementations of ESA: (1) leveraging the MultiheadAttention module from PyTorch, (2) leveraging the MultiHeadDispatch module from xFormers, (3) a manual implementation of multihead attention, relying on PyTorch’s scaled_dot_product_attention function, (4) a manual implementation of multihead attention, relying on xFormers’ memory_efficient_attention, and (5) a naive implementation. Options (1) - (4) can all make use of efficient and fast implementations. However, we have observed performance differences between the 4 implementations as well as compared to a naive implementation. This behaviour is likely due to the different low-level kernel implementations. Moreover, Flash attention does not currently support custom attention masks, as there is little interest for such functionality from a language modelling perspective.
@@
Although masks can be computed efficiently during training, all frameworks require the last two dimensions of the input mask tensor to be of shape $(L_{n},L_{n})$ for nodes or $(L_{e},L_{e})$ for edges, effectively squaring the number of nodes or edges. However, the mask tensors are sparse and a sparse tensor alternative could greatly reduce the memory consumption for large and dense graphs. Such an option exists for native PyTorch attention, but it is currently broken. Additionally, even a single boolean requires an entire byte of storage in PyTorch, increasing the memory usage by 8 times compared to the theoretical requirement of 1 bit.
@@
Another possible optimisation would be to use nested (ragged) tensors to represent graphs, since padding is currently necessary to ensure identical dimensions for attention. A prototype nested tensor attention is available in PyTorch; however, not all the required operations are supported, and converting between normal and nested tensors is slow.
@@
For all implementations, it is required that the mask tensor is repeated by the number of attention heads (e.g. 8 or 16). However, a notable bottleneck is encountered for the MultiheadAttention and MultiHeadDispatch variants described above, which require that the repeats occur in the batch dimension, i.e. requiring 3D mask tensors of shape $(B\times H,L,L)$, where $H$ is the number of heads. The other two efficient implementations require a 4D mask instead, i.e. $(B,H,L,L)$, where one can use PyTorch’s expand function instead of repeat. The expand alternative does not use any additional memory, while repeat requires ×$\times H$ memory. Note that it is not possible to reshape the 4D tensor created using expand without using additional memory.
@@
Finally, we noticed a limitation involving PyTorch’s nonzero() tensor method, which is required as part of the edge masking algorithm (Introduction). This is covered in detail in SI 12. Currently, the nonzero() method fails for very dense graphs. A fix would require an update to 64-bit integer limits.
Supplementary Information SI 8 Helper functions
@@
consecutive is a helper function that generates consecutive numbers starting from 0, with a length specified in its tensor argument as the difference between adjacent elements, and a second integer argument used for the last length computation, e.g. consecutive($[1,4,6]$, $10$) $=[0,1,2,0,1,0,1,2,3]$, and first_unique_index finds the first occurrence of each unique element in the tensor (sorted), e.g. first_unique_index($[3,2,3,4,2]$) $=[1,0,3]$. The implementations are available in our code base.
Supplementary Information SI 9 Experimental setup
@@
We follow a simple and universal experimental protocol to ensure that it is possible to compare the results of different methods and to evaluate a large number of datasets with high throughput. In Section SI 9.1 we describe the grid search approach and the search parameters used. For basic hyperparameters such as batch size and learning rate, we chose a number of reasonable hyperparameters and settings for all methods, regardless of their nature (GNN or attention-based). This includes the AdamW optimiser [loshchilov2018decoupled], more specifically the 8-bit version [dettmers2022bit], learning rate ($0.0001$), batch size ($128$), mixed precision training with the bfloat16 tensor format, early stopping with a patience of 30 epochs (100 for very small datasets such as freesolv), and gradient clipping (set to the default value of $0.5$). Furthermore, we used a simple learning rate scheduler that halved the learning rate if no improvement was encountered for 15 epochs (half the early stopping patience). If these parameters led to out-of-memory errors, we attempted reducing the batch size by 2 until the error was fixed, or reducing the hidden dimension as a last resort.
@@
For a fair and comprehensive evaluation, we tune each algorithm for each dataset using a grid search approach and a selection of reasonable hyperparameters. These include the number of layers, the number of attention heads for GAT(v2) and graph transformers, dropout, and hidden dimensions. For graph-level GNNs, we evaluated configurations with 4 to 6 layers, hidden dimensions in $\{128,256,512\}$, the number of GAT heads in $\{8,16\}$, and GAT dropout in $\{0,0.2\}$. For node-level GNNs, we adapted the search since these tasks are more sensitive to overfitting, and used a number of layers in $\{1,2,4,6\}$, hidden dimensions in $\{64,128,256\}$, the same GAT heads and dropout settings, and dropout after each GNN layer in $\{0,0.2\}$. For Graphormer, TokenGT, and GraphGPS, we evaluated models with the number of layers in $\{4,6,8,10\}$, the number of attention heads in $\{4,8,16\}$, and hidden dimensions in $\{128,256,512\}$. For ESA, we generally focused on models with 6 to 10 layers, with SABs at the start and end and MABs in the middle. The hidden dimensions are selected from $\{256,512\}$ and the number of attention heads from $\{8,16,32\}$. We evaluated pre-LN and post-LN architectures, standard and gated MLPs, and different MLP hidden dimensions and number of MLP layers on a dataset-by-dataset basis. The best configuration is selected based on the validation loss, and results are reported on the test set from 5 different runs. Based on recent reports on the performance of GNNs [luo2024classic], we augmented all 6 GNN baselines with residual connections and normalisation layers for each graph convolutional layer. These strategies are not part of the original message passing specification, but lead to substantial uplifts. For datasets with established splits, such as cifar10, mnist, or zinc, we use the available splits. For dockstring, only train and test splits are available, so we randomly extract 20,000 train molecules for validation. Otherwise, we generate our own splits with a train/validation/test ratio of 80%/10%/10%. All models are trained and evaluated using mixed-precision training with the bfloat16 tensor format.
SI 9.2 Metrics
@@
Given the scale of our evaluation, it is crucial to use an appropriate selection of performance metrics. To this end, we selected the metrics according to established and recent literature. For classification tasks, there is a growing consensus that Matthew’s correlation coefficient (MCC) is the preferred metric over alternatives such as accuracy, F-score, and the area under the receiver operating characteristic curve (AUROC or ROC-AUC) [Chicco2020, 9440903, Chicco2021, STOICA2024109511]. The AUROC in particular has been shown to be problematic [Chicco2023, 10.1145/1557019.1557021, 10.1145/3593013.3594100]. The MCC is an informative measure of a classifier’s performance as it summarises the four basic rates of a confusion matrix: sensitivity, specificity, precision, and negative predictive value [Chicco2023]. Similarly, the $\text{R}^{2}$ has been proven to be more informative than alternatives such as the mean absolute or squared errors for regression tasks [Chicco2021-bg]. Thus, our first choices for reporting results are the MCC and $\text{R}^{2}$, depending on the task. For comparison with leaderboard results and for specialised fields such as quantum mechanics, we also report comparable metrics (i.e. accuracy, mean absolute error, or root mean squared error).
SI 9.3 Baselines
@@
We include classic message passing baselines in the form of GCN, GAT, and GIN due to their recent resurgence against sophisticated graph transformers and their widespread use. We also include the improved GATv2 [brody2022how] to complement GAT, and PNA for being neglected in other works despite its remarkable empirical performance. We complete the message passing baselines by including DropGNN [wang2023mathscrnwl], a family of provably expressive GNNs that can solve tasks beyond 1-WL in an efficient manner by randomly dropping nodes. As in the original paper, we use GIN as the main underlying mechanism and label this technique DropGIN. Regarding transformer baselines, we select Graphormer, TokenGT, and GraphGPS, not only because of their widespread use but also because of generally outperforming previous generation graph transformers such as SAN. This selection is balanced in the sense that Graphormer and TokenGT are part of a class of algorithms that focuses on representing the graph structure through encodings and token identifiers, while GraphGPS relies on GNNs and is thus a hybrid approach.
Supplementary Information SI 10 Transfer learning setup
The transfer learning setup consists of randomly selected training, validation, and test sets of 25K, 5K, and, respectively, 10K molecules with GW calculations (from the total of 133,885). This setup mimics the low amounts of high quality/fidelity data available in drug discovery and quantum simulation projects. In the transductive case, the entire dataset with DFT targets is used for pre-training (including the 10K test set compounds, but only with DFT-level measurements), while in the inductive setting the 10K set is completely excluded. Here, we perform transfer learning by pre-training a model on the DFT target for a fixed number of epochs (150) and then fine-tuning it on the subset of 25K GW calculations. In the transductive case, pre-training occurs on the full set of 133K DFT calculations, while in the inductive case the DFT test set values are removed (note that the evaluation is done on the test set GW measurements).
Supplementary Information SI 11 Infected graph generation
@@
The infected graphs are generated using the InfectionDataset from PyTorch Geometric. We generate two Erdős–Rényi (ER) graphs with different sizes:
@@
Supplementary Information SI 12 Adaptations for Open Catalyst Project
@@
Extending a given model to work with 3D data is not trivial, as demonstrated by the follow-up paper dedicated to extending and benchmarking Graphormer on 3D molecular problems [shi2023benchmarking]. As described in that paper, for the Open Catalyst Project (OCP) data, certain pre-processing steps are taken to ensure satisfactory performance. Concretely, a set of Gaussian basis functions is used to encode atomic distances, which are not used in their raw form. The idea of encoding raw quantities to achieve expressive and orthogonal representations has also been studied by Gasteiger et al. for DimeNet [gasteiger_dimenet_2020], taking things further and using Bessel functions. This idea is prevalent in the literature and shows some of the complications of working with 3D coordinates.